# rQUFguide.bib

### v1

@@ 4 A Simple Model
Clearly in Section 3.5 we see that the events are endogenously excited by themselves. This stylized fact, along with the general applicability and versatility of the Hawkes Process as discussed in Section 2.3, is the main motivation of using a self exciting point process as our LOB model. We propose a model of the LOB dynamics where the events are driven through a multidimensional Hawkes Process and there exist some state variables (such as the top of the book volume at Ask) of the LOB which evolve along with the point process. As discussed in [jain2023hawkes], a Hawkes model alone is unable to capture dynamics of a small tick order book. The assumptions of the model are as follows.
  • Firstly, the LOB is dense (i.e. not sparse). There are no empty levels between the deep levels and the top level. This assumption clearly fails for small tick LOBs as we have shown in Section 3.4.
  • Secondly, the price improvements, i.e. in-spread orders, happen uniformly at a 1 tick distance from the top of the book. This is clearly not the case in small tick stocks since price improvements happen in multiple ticks.
  • Thirdly, the top of the book contains most of the liquidity of the LOB. As shown in Section 3.3, for small and some medium tick stocks, the maximum liquidity actually lies much deeper in the book.
  • Finally, the assumption that events in the price levels other than the top two have no contribution to the LOB dynamics is too strong. Section 3.5 depicts the same.
@@
Therefore in order to enhance this model and to relax these assumptions, we propose to create a model of the LOB which not only captures the top of the book dynamics but also is able to model the deeper price levels’ events. Simply increasing the dimensions of the Hawkes Process model to incorporate events in the deeper LOB, and to incorporate in-spread orders at all levels from top to the opposite side’s top, is infeasible. This is because the Hawkes Process calibration becomes exponentially harder as we increase the dimensions of the process. Even if we could calibrate the process, the question of model parsimony becomes important. Therefore we propose here to model events happening in a range of price levels as a single dimension in the Hawkes Process. The range of these ’fat queues’ is denoted by their ’width’. These widths are dynamic quantities which evolve with the events in the LOB. These width variables along with the queue sizes are what we call ’state variables’ of the LOB. We describe these state variables below and then propose a simple model where the parameters of the Hawkes Process and the LOB State Variables are independently calibrated.
4.1 Formulation:
@@
From Section 3.5, it is clear that for all stocks, there needs to be a demarkation between the top of the book and the deeper levels, for both LOs and COs. This, along with LOs at the in-spread levels, and MOs, constitute all the dimensions of the book. The choice of how deep we would like to model the LOB is a pertinent question. As a heuristic, we choose the 50th percentile of the average shape of the LOB as our maximum depth that we would like to model. We define this quantity as $M_{0.5}$ which is a constant in time.
@@
$m^{(\zeta)}_{D}(t)$ is defined as the width of the deeper levels. This is a random process which evolves dynamically with the LOB events. We initiate this using $m^{(\zeta)}_{D}(0)=M_{0.5}-0.5s(0)-m^{(\zeta)}_{T}(0)$.
For example, in Figure 10, the second best ask quote is at 1133.74 \$ and mid-price $+M_{0.5}$ is at 1133.88 \$. Therefore $m^{(ask)}_{D}(t)=14$ ticks. This random process has a constraint that we can never have more than $M_{0.5}+1$ levels on each side of the book in the model. This choice is motivated by model parsimony as well as the hypothesis that any quotes deeper than $M_{0.5}$ ticks away from the mid-price have neglible contribution to the LOB’s top of the book dynamics. Mathematically, this constraint can be written as $M_{0.5}-0.5s(t)-m^{(\zeta)}_{T}(t)-m^{(\zeta)}_{D}(t)>=-1$ and $m^{(\zeta)}_{D}(t)>=1$. Therefore it is possible that while the LOB is evolving, the number of levels modelled (on each side) by these three fat-queues is much less than $M_{0.5}$. We discuss in the following how we maintain these constraints.
There are therefore three wide queues, at each side, we keep track of in this model - the in-spread queue (i.e. for Limit Orders arriving inside the spread), the top queue and the deeper queue. Each of these queues have a queue size which is the total order volume resting in the queue at that instant of time. Naturally, the queue size for in-spread queues is always zero. We denote the queue size at the top queue and the deeper queue by $Q^{(\zeta)}_{T}(t),Q^{(\zeta)}_{D}(t)$ respectively. These are again random processes with domain $\mathbb{N}$. There are total 12 events that can happen in these queues which are, on each side of the book, LOs in the three wide queues, COs in the two queues with non-zero queue size, and MOs at the top of the book. The set of events is:
:=$\mathcal{E}:=\{LO_{\text{ask}_{D}},CO_{\text{ask}_{D}},LO_{\text{ask}_{T}},CO_{\text{ask}_{T}},MO_{\text{ask}_{T}},LO_{\text{ask}_{IS}},LO_{\text{bid}_{IS}},LO_{\text{bid}_{T}},CO_{\text{bid}_{T}},MO_{\text{bid}_{T}},LO_{\text{bid}_{D}},CO_{\text{bid}_{D}}\}$
4.2 Methodology
Since the order book queues are not restricted to a single tick width, determining the precise price level at which a simulated order is placed at becomes a critical consideration. The Hawkes process model, incorporating the 12 aforementioned event types, provides the event time and type upon simulation, but does not specify the exact price associated with an order—particularly for LOS and COs. To address this, we derive difference equations for the LOB state variables corresponding to each event type. The subsequent section details the event types and their impact on this coarse-grained LOB representation.
@@ 4.2 Methodology
    $LO_{\zeta_{IS}}$:
    Since IS orders can arrive at any price point between the top of the book and the top of the opposing side, it is necessary to establish a methodology for determining the precise price point at which such an order will arrive. Let $\eta_{IS}\in\mathbb{N}$ be defined as the number of ticks from the top at side $\zeta$ where this new in-spread order arrives at. We sample this $\eta_{IS}$ from a calibrated stationary distribution $\Pi(\hat{\eta}_{IS})$ where $\hat{\eta}_{IS}$ is a calibrated parameter. We denote this relation as $\eta_{IS}(\in\mathbb{N})\sim\Pi(\hat{\eta}_{IS})$ from here onwards. $\eta_{IS}$ is capped at the current spread in ticks and floored by one tick. The choice of distribution shape and the calibration of the parameter $\hat{\eta}_{IS}$ is directly related to the stylized fact of sparsity showcased in section 3.4.
    Also, the order size of this new order $\kappa_{IS}\in\mathbb{N}$ is sampled from a calibrated stationary distribution $\Pi(\hat{\kappa}_{IS})$. This distribution of the order size can contain spikes at round numbers similar to the empirical observations found by [jain2023hawkes].
    The following changes happen to the five random processes mentioned above:
    $\eta_{IS}$ $\sim\Pi(\hat{\eta}_{IS})$ (1)
@@
    Since this event is an event which increases $m^{(\zeta)}_{D}(t)$, we need to check the constraint mentioned in Definition 3. If the constraint is violated, we do the following purging of the width of the deeper levels to maintain the constraint of modelling at most the top $M_{0.5}$ levels in the LOB.
    $m^{(\zeta)}_{D}(t^{+})$ $=M_{0.5}-0.5s(t^{+})-m^{(\zeta)}_{T}(t^{+})$ (8)
@@
    Here $\Delta Q^{(\zeta)}_{D}(t)$ is defined as a partition of the existing $Q^{(\zeta)}_{D}(t)$. This partition directly depends on the number of levels in the partition ($m^{(\zeta)}_{D}(t^{+})-m^{(\zeta)}_{T}(t)$) and the number of levels in the parent ($m^{(\zeta)}_{D}(t)$). Figure 11 shows the same logic pictorially. In this instance, a new in-spread order arrives at the price point of 1133.66 \$, corresponding to $\eta_{IS}=6$ ticks. According to equation (2), the mid-price jump $m^{(\zeta)}_{T}(t^{+})$ increases to 6 ticks from the previous 2 ticks. However, the deep width variable $m^{(\zeta)}_{D}(t)$ cannot be further increased, as doing so would exceed the modeled levels limit (indicated by the dashed red line). Consequently, the orders at the price level of 1133.94 \$ are purged, ensuring that $m^{(\zeta)}_{D}(t^{+})$ no longer violates this constraint. This example is derived from a real Limit Order Book (LOB) in our dataset. In the Hawkes process simulations, this level of fine-grained information in the deep dimension is unavailable. Therefore, it becomes necessary to purge using a partition function based on an assumed liquidity distribution at the deeper levels. This partition function can be calibrated according to the shape of the order book, as outlined in section 3.3.
@@
    $LO_{\zeta_{T}}$:
    Since the top of the book consists of multiple price levels, we need to follow the same strategy as above to define which price level the LO will arrive at. Let $\eta_{T}(\in\mathbb{N})\sim\Pi(\hat{\eta}_{T})$ be defined as the number of ticks from the top at side $\zeta$ where this new order arrives at. $\eta_{T}$ is capped at $m^{(\zeta)}_{T}(t)$ and floored at zero. Again, the calibration of the distribution shape and the parameter can be informed using the sparsity stylized facts measured in section 3.4. Finally, $\kappa_{T}(\in\mathbb{N})\sim\Pi(\hat{\kappa}_{T})$ is the order size of this LO. If $\eta_{T}$ is 0, the LO is simply added to the existing queue at the top and no changes occur in the 5 processes except $Q^{(\zeta)}_{T}(t^{+})=Q^{(\zeta)}_{T}(t)+\kappa_{T}$. Otherwise, the following changes happen to the random processes mentioned above:
    $\eta_{T}$ $\sim\Pi(\hat{\eta}_{T})$ (11)
@@
    $CO_{\zeta_{T}}$:
    If the CO does not lead to a queue depletion, i.e. if the number of orders at the top is more than one, then we randomly select some quantity from the existing queue at the top and cancel that quantity. The top of the book queue size changes as follows $Q^{(\zeta)}_{T}(t^{+})=Q^{(\zeta)}_{T}(t)-\kappa_{CO}$ where $\kappa_{CO}$ is the randomly selected quantity. If the CO leads to a queue depletion (QD), we need to calculate the new top of the book width as well as deplete the deeper fat queue by this new top level’s total volume. Assuming a sparse structure within the deep fat queue (informed by section 3.4), we sample a random new top width denoted by $\eta_{T+1}$. To calculate the total volume at this new level, we make use of the partition function mentioned earlier to deplete the deep fat queue. Mathematically, we perform the following to redefine a new top of the book.
@@
    Here, $\eta_{T+1}(\in\mathbb{N})\sim\Pi(\hat{\eta}_{T+1})$ is defined as the number of levels the new top has after the queue depletion event.
@@
    $MO_{\zeta}$:
    Market Orders, mathematically, behave exactly the same as Cancel Orders. If the size of the market order $\kappa_{MO}(\in\mathbb{N})\sim\Pi(\hat{\kappa}_{MO})$ is less than the existing queue size, there is a simple reduction in volume at the top, according to time priority, and we report a trade. The top of the book queue size changes as follows $Q^{(\zeta)}_{T}(t^{+})=Q^{(\zeta)}_{T}(t)-\kappa_{MO}$. If the MO leads to a queue depletion, the same logical waterfall as the one above is followed. If a market order depletes more than one price level, we repeat the process until we completely execute the MO quantity. Thus a MO can deplete multiple price levels using this methodology. This clearly means that the stylized facts related to the price moves induced by trades 3.2 is modelled implicitly.
@@
    $LO_{\zeta_{D}}$:
    Since we do not keep track of the fine grained orders or price levels in the deeper dimension, we simply increment the total volume present there with this new LO size $\kappa_{D}(\in\mathbb{N})\sim\Pi(\hat{\kappa}_{D})$.
@@
    $CO_{\zeta_{D}}$:
    In case of a QD event, if $m^{(\zeta)}_{D}(t)+m^{(\zeta)}_{T}(t)+0.5s(t)\leq M_{0.5}$, i.e. the width of the deep fat queue is not violating the constraint in Defintion 3, we absorb the new empty levels in $m^{(\zeta)}_{T}(t)$ and create a new deep fat queue with the following width and volume:
    $m^{(\zeta)}_{D}(t^{+})$ $=M_{0.5}-m^{(\zeta)}_{T}(t)-0.5s(t)$ (21)
    $Q^{(\zeta)}_{D}(t^{+})$ $\sim\Pi(\hat{Q}_{D},m^{(\zeta)}_{D}(t^{+}))$ (22)
@@
    If the constraint is violated, we set $m^{(\zeta)}_{D}(t^{+})$ to 1, sample $Q^{(\zeta)}_{D}(t^{+})$ from $\Pi(\hat{Q}_{D},m^{(\zeta)}_{D}(t^{+}))$ and let the deeper queue have just one non empty level since we cannot go deeper than $M_{0.5}+1$ in total depth of the LOB.
The 12 events form a mutually exciting 12 dimensional Hawkes Process with their intensities as follows from [jain2023hawkes]. For $i=1\ldots 12,d=12$, let a $d$-dimensional Hawkes process have as the intensity of the process $\lambda^{(i)}(t)$ and the associated counting process $N^{(i)}(t)$ for $i=1,\ldots,d$. Here, $\mu^{(i)}$ is the exogenous intensity of the $i$-th dimension and $\phi^{(j\rightarrow i)}(t)$ is the excitation term from $j$-th dimension to $i$-th dimension. $\alpha$ is defined as the scale parameter for the in-spread intensity and $\beta$ is defined as the shape parameter.
@@
4.3 Calibration:
Since the Hawkes Dynamics is assumed to be independent of the LOB state variables, and since the distributions of the order sizes, order distances from top and queue sizes are unconditional on any state variable and stationary, we can decouple the calibration to a parametric estimation of the distributions and a non-parametric estimation of the Hawkes kernels. We suggest the methodology described in [jain2023hawkes], combining [kirchner2017estimation] and [bacry2016estimation], for the latter. The authors also provide visualization of the estimated kernel norms for INTC (Large Tick), AAPL (Medium Tick), TSLA and AMZN (Small Tick). For the former, we suggest making use of the Maximum Likelihood method with a prior that all these distributions are Geometric distributions. In Figure 12, we showcase the three $\eta$s distributions (note the shape looks exactly like the Geometric distribution) and their calibrated parameter: $\hat{\eta}_{(.)}$ across three stocks: INTC (large tick), AAPL (medium tick) and GOOG (small tick). Optionally, following [jain2023hawkes, LuAbergel2018], one can add spikes in the probability distribution function to account for the traders’ preference of round numbers in the order sizes and queue sizes. For the partition functions, in the following section, we make use of a simple linear partition function - i.e. we assume that the shape of the book in the deeper LOB is flat. This is clearly a strong assumption (Section 3.2), and a better choice could be using the average shape of the book instead.
(a) $\hat{\eta}_{IS}$ Calibration
(b) $\hat{\eta}_{T}$ Calibration
(c) $\hat{\eta}_{T+1}$ Calibration
Figure 12: $\hat{\eta}_{(.)}$ Calibration: $\hat{\eta}_{(.)}$ is the parameter which controls the sparsity of the LOB. Here we show three examples, one from each category of stocks, of empirical histograms of the sparsity levels. On a log-log scale it can be clearly seen that the distribution is similar to the geometric distribution. Inset we also show the calibrated parameters with this prior for all three stocks - we observe the calibrated parameters move from near one to near zero as we decrease the rel. tick size. This indicates an increase in sparsity with a decrease in rel. tick size.
5 Results
5.1 Stylized Facts:
In this section we describe the dynamics of the stylized facts defined above in Section 3 from the above model formulation.
5.1.1 Bid-Ask Spread and Price Moves:
It is easy to see that the spread process is decreased by $\eta_{IS}$ whenever an in-spread LO occurs and increased by $m^{(\zeta)}_{T}(t)$ whenever a queue depletion event (QD) occurs on side $\zeta$. Therefore we can write the differential equation of spread as:

### v2

@@ 3 A Wide-Queue Hawkes Model for Limit Order Books
Clearly in Section 2.5 we see that the events are endogenously excited by themselves. This stylized fact, along with the general applicability and versatility of the Hawkes Process as discussed in Section 1.2, is the main motivation of using a self exciting point process as our LOB model. We propose a model of the LOB dynamics where the events are driven through a multidimensional Hawkes Process and there exist some state variables (such as the top of the book volume) of the LOB which evolve along with the point process. As discussed in [jain2023hawkes], a Hawkes model alone is unable to capture dynamics of a small tick order book. The assumptions of the model in [jain2023hawkes] are as follows. Firstly, the LOB is dense (i.e. not sparse). There are no empty levels between the deep levels and the top level. This assumption clearly fails for small tick LOBs as we have shown in Section 2.4. Secondly, the price improvements, i.e. in-spread orders, happen uniformly at a one tick distance from the top of the book. This is clearly not the case in small tick stocks since price improvements happen in multiple ticks (Section 2.2). Thirdly, the top of the book contains most of the liquidity of the LOB. As shown in Section 2.3, for small and some medium tick stocks, the maximum liquidity actually lies much deeper in the book. Finally, the assumption that events in the price levels other than the top two have no contribution to the LOB dynamics is too strong. Section 2.5 depicts the same.
@@
Therefore in order to enhance this model and to relax previous assumptions, we propose to create a model of the LOB which not only captures the top of the book dynamics but also is able to model the deeper price levels’ events. Simply increasing the dimensions of the Hawkes Process model to incorporate events in the deeper LOB, and to incorporate in-spread orders at all levels from top to the opposite side’s top, is infeasible. This is because the Hawkes Process calibration becomes exponentially harder as we increase the dimensions of the process. Even if we could calibrate the process, the question of model parsimony becomes important. Further, we conclude from Section 2.5 that in-spread orders should be modelled differently for different types of stocks however there is uniformity in other order types’ behaviour. Additionally, the top of the LOB is a very special price level and the events happening there are excited very differently from the rest of the LOB. The deeper levels seem to behave more or less in a similar manner. Therefore we propose here to model events happening in a range of price levels as a single dimension in the Hawkes Process. The range of these ‘wide queues’ is denoted by their ‘width’. These widths are dynamic quantities which evolve with the events in the LOB. These width variables along with the queue sizes are what we call ‘state variables’ of the LOB. We describe these state variables below and then propose a simple model where the parameters of the Hawkes Process and the LOB State Variables are independently calibrated.
3.1 Model Formulation:
@@
From Section 2.5, it is clear that for all stocks, there needs to be a demarkation between the top of the book and the deeper levels, for both LOs and COs. This, along with LOs at the in-spread levels, and MOs, constitute all the dimensions of the book. The choice of how deep we would like to model the LOB is a pertinent question. As a heuristic, we choose the 50th percentile of the average shape of the LOB as our maximum depth that we would like to model. We define this quantity as $M_{\frac{1}{2}}$ which is a constant in time for each stock. A useful reference for the exact value of $M_{\frac{1}{2}}$ is given as the 2nd Quartile in Table 2.
@@
$m^{(\zeta)}_{D}(t)$ is defined as the width of the deeper levels. This is a random process which evolves dynamically with the LOB events. We initiate this using $m^{(\zeta)}_{D}(0)=M_{\frac{1}{2}}-\frac{1}{2}s(0)-m^{(\zeta)}_{T}(0)$.
For example, in Figure 10, the second best ask quote is at 1133.74 \$ and mid-price $+M_{\frac{1}{2}}$ is at 1133.88 \$. Therefore $m^{(ask)}_{D}(t)=14$ ticks. This random process has a constraint that we can never have more than $M_{\frac{1}{2}}+1$ levels on each side of the book in the model. This choice is motivated by model parsimony as well as the hypothesis that any quotes deeper than $M_{\frac{1}{2}}$ ticks away from the mid-price have neglible contribution to the LOB’s top of the book dynamics. Mathematically, this constraint can be written as $M_{\frac{1}{2}}-\frac{1}{2}s(t)-m^{(\zeta)}_{T}(t)-m^{(\zeta)}_{D}(t)>=-1$ and $m^{(\zeta)}_{D}(t)>=1$. Therefore it is possible that while the LOB is evolving, the number of levels modelled (on each side) by these three fat-queues is much less than $M_{\frac{1}{2}}$. We discuss in the following how we maintain these constraints.
There are therefore two wide queues, at each side, we keep track of in this model - the top queue and the deeper queue. Each of these queues have a queue size which is the total order volume resting in the queue at that instant of time. We denote the queue size at the top queue and the deeper queue by $Q^{(\zeta)}_{T}(t),Q^{(\zeta)}_{D}(t)$ respectively. For best price at side $\zeta$ denoted by $p_{\zeta}(t)$, they are defined as the following:
$Q^{(\zeta)}_{T}(t)$ $:=\sum^{m^{(\zeta)}_{T}(t)-1}_{x=0}Q(p_{\zeta}(t)+\delta x,t)$ (10)
$Q^{(\zeta)}_{D}(t)$ $:=\sum^{m^{(\zeta)}_{D}(t)+m^{(\zeta)}_{T}(t)}_{x=m^{(\zeta)}_{T}(t)}Q(p_{\zeta}(t)+\delta x,t)$ (11)
These are again random processes with domain $\mathbb{N}$. There are total 12 events that can happen in these queues which are, on each side of the book, LOs in the two wide queues, COs in the two queues, In-Spread LOs (IS), and MOs at the top of the book. The set of events is:
$\mathcal{E}:=$$\{LO_{\text{ask}_{D}},CO_{\text{ask}_{D}},LO_{\text{ask}_{T}},CO_{\text{ask}_{T}},MO_{\text{ask}_{T}},LO_{\text{ask}_{IS}},$
$LO_{\text{bid}_{IS}},LO_{\text{bid}_{T}},CO_{\text{bid}_{T}},MO_{\text{bid}_{T}},LO_{\text{bid}_{D}},CO_{\text{bid}_{D}}\}$
3.2 Events and State Variables Dynamics
Since the order book queues are not restricted to a single tick width, determining the precise price level at which a simulated order is placed at becomes a critical consideration. The Hawkes process model, incorporating the 12 aforementioned event types, provides the event time and type upon simulation, but does not specify the exact price associated with an order—particularly for LOs and COs. To address this, we derive difference equations for the LOB state variables corresponding to each event type. This section details the event types and their impact on this coarse-grained LOB representation.
@@ 3.2 Events and State Variables Dynamics
    In-Spread Limit Orders - $LO_{\zeta_{IS}}$:
    Since IS orders can arrive at any price point between the top of the book and the top of the opposing side, it is necessary to establish a methodology for determining the precise price point at which such an order will arrive. Let $\eta_{IS}\in\mathbb{N}$ be defined as the number of ticks from the top at side $\zeta$ where this new in-spread order (of side $\zeta$ i.e. $LO_{\zeta_{IS}}\in\mathcal{E}$) arrives at. We sample this $\eta_{IS}$ from a calibrated stationary distribution $\Pi_{\eta}(\hat{\eta}_{IS})$ [^3] A note on notation: We use $\Pi_{a}(\hat{a}_{e})$ as the notation for a calibrated stationary, parametric probability distribution function with parameter $\hat{a}_{e}$ for the random variable $a_{e}$ affecting the event $e$’s dynamics. Note that the distribution choice is consistent across all events for a random variable type (for example $\eta_{IS/T/T+1}$ all share the same class of distribution in the model) and therefore we use the notation of $\Pi_{a}(.)$. Each distribution itself can differ in the family of distributions it belongs to. For example, one might choose, for a Large Tick asset, the Dirac Delta distribution for $\Pi_{\eta}(\hat{\eta}_{IS})$ and the Geometric distribution for order sizes $\Pi_{\kappa}(\hat{\kappa}_{IS})$. where $\hat{\eta}_{IS}$ is a calibrated parameter. We denote this relation as $\eta_{IS}(\in\mathbb{N})\sim\Pi_{\eta}(\hat{\eta}_{IS})$ from here onwards. $\eta_{IS}$ is capped at the current spread in ticks and floored by one tick. The choice of distribution shape and the calibration of the parameter $\hat{\eta}_{IS}$ is directly related to the stylized fact of sparsity showcased in section 2.4. For example, in the following section, we utilize the Geometric Distribution for $\Pi_{\eta}(.)$.
    Also, the order size of this new order $\kappa_{IS}\in\mathbb{N}$ is sampled from a calibrated stationary distribution $\Pi_{\kappa}(\hat{\kappa}_{IS})$. This distribution of the order size can contain spikes at round numbers similar to the empirical observations found by [jain2023hawkes].
    The following changes happen to the five random processes mentioned above, as well as the mid-price $p_{mid}(t)$:
    $\eta_{IS}$ $\sim\Pi_{\eta}(\hat{\eta}_{IS})$ (12)
    $\kappa_{IS}$ $\sim\Pi_{\kappa}(\hat{\kappa}_{IS})$ (13)
@@
    Since this event is an event which increases $m^{(\zeta)}_{D}(t)$, we need to check the constraint mentioned in Definition 3. If the constraint is violated, we do the following purging of the width of the deeper levels to maintain the constraint of modelling at most the top $M_{\frac{1}{2}}$ levels in the LOB.
    $m^{(\zeta)}_{D}(t^{+})$ $=M_{\frac{1}{2}}-\frac{1}{2}s(t^{+})-m^{(\zeta)}_{T}(t^{+})$ (20)
@@
    Here $\Delta Q^{(\zeta)}_{D}(t)$ is defined as a partition of the existing $Q^{(\zeta)}_{D}(t)$. This partition directly depends on the number of levels in the partition ($m^{(\zeta)}_{D}(t^{+})-m^{(\zeta)}_{T}(t)$) and the number of levels in the parent ($m^{(\zeta)}_{D}(t)$). Figure 11 shows the same logic pictorially. In this instance, a new in-spread order arrives at the price point of 1133.66 \$, corresponding to $\eta_{IS}=6$ ticks. According to Equation 14, $m^{(\zeta)}_{T}(t^{+})$ increases to 6 ticks from the previous 2 ticks. However, the deep width variable $m^{(\zeta)}_{D}(t)$ cannot be further increased, as doing so would exceed the modeled levels limit (indicated by the dashed red line). Consequently, the orders at the price level of 1133.94 \$ are purged, ensuring that $m^{(\zeta)}_{D}(t^{+})$ no longer violates this constraint. This example is derived from a real Limit Order Book (LOB) in our dataset. In the Hawkes process simulations, this level of fine-grained information in the deep dimension is unavailable. Therefore, it becomes necessary to purge using a partition function based on an assumed liquidity distribution at the deeper levels. This partition function can be calibrated according to the shape of the order book, as outlined in Section 2.3.
@@
    Limit Orders at Top - $LO_{\zeta_{T}}$:
    Since the top of the book consists of multiple price levels, we need to follow the same strategy as above to define which price level the LO will arrive at. Let $\eta_{T}(\in\mathbb{N})\sim\Pi_{\eta}(\hat{\eta}_{T})$ be defined as the number of ticks from the top at side $\zeta$ where this new order arrives at. $\eta_{T}$ is capped at $m^{(\zeta)}_{T}(t)$ and floored at zero. Again, the calibration of the distribution shape and the parameter can be informed using the sparsity stylized facts measured in section 2.4. Finally, $\kappa_{T}(\in\mathbb{N})\sim\Pi_{\eta}(\hat{\kappa}_{T})$ is the order size of this LO. If $\eta_{T}$ is 0, the LO is simply added to the existing queue at the top and no changes occur in the 5 processes except $Q^{(\zeta)}_{T}(t^{+})=Q^{(\zeta)}_{T}(t)+\kappa_{T}$. Otherwise, the following changes happen to the random processes mentioned above:
    $\eta_{T}$ $\sim\Pi_{\eta}(\hat{\eta}_{T})$ (23)
@@
    Cancel Orders at Top - $CO_{\zeta_{T}}$:
    If the CO does not lead to a queue depletion, i.e. if the number of orders at the top is more than one, then we randomly select some quantity from the existing queue at the top and cancel that quantity. The top of the book queue size changes as follows $Q^{(\zeta)}_{T}(t^{+})=Q^{(\zeta)}_{T}(t)-\kappa_{CO}$ where $\kappa_{CO}$ is the randomly selected quantity. If the CO leads to a queue depletion (QD), we need to calculate the new top of the book width as well as deplete the deeper wide queue by this new top level’s total volume. Assuming a sparse structure within the deep wide queue (informed by section 2.4), we sample a random new top width denoted by $\eta_{T+1}$. Here, $\eta_{T+1}(\in\mathbb{N})\sim\Pi_{\eta}(\hat{\eta}_{T+1})$ is defined as the number of levels the new top has after the queue depletion event. To calculate the total volume at this new level, we make use of the partition function mentioned earlier to deplete the deep wide queue. Mathematically, we perform the following to redefine a new top of the book.
    $\eta_{T+1}$ $\sim\Pi_{\eta}(\hat{\eta}_{T+1})$ (28)
@@
    $s(t^{+})$ $=s(t)+m^{(\zeta)}_{T}(t)$ (33)
    $p_{mid}(t^{+})$ $=p_{mid}(t)+z\frac{m^{(\zeta)}_{T}(t)}{2}$ (34)
@@
    Market Orders - $MO_{\zeta}$:
    Market Orders, mathematically, behave exactly the same as Cancel Orders. If the size of the market order $\kappa_{MO}(\in\mathbb{N})\sim\Pi_{\kappa}(\hat{\kappa}_{MO})$ is less than the existing queue size, there is a simple reduction in volume at the top, according to time priority, and we report a trade. The top of the book queue size changes as follows $Q^{(\zeta)}_{T}(t^{+})=Q^{(\zeta)}_{T}(t)-\kappa_{MO}$. If the MO leads to a queue depletion, the same logical waterfall as the one above is followed. If a market order depletes more than one price level, we repeat the process until we completely execute the MO quantity. Thus a MO can deplete multiple price levels using this methodology. This clearly means that the stylized facts related to the price moves induced by Section 2.2 is modelled implicitly.
@@
    Deep Limit Orders - $LO_{\zeta_{D}}$:
    Since we do not keep track of the fine grained orders or price levels in the deeper dimension, we simply increment the total volume present there with this new LO size $\kappa_{D}(\in\mathbb{N})\sim\Pi_{\kappa}(\hat{\kappa}_{D})$.
@@
    Deep Cancel Orders - $CO_{\zeta_{D}}$:
    Again, if there is no queue-depletion event, we simply remove the CO volume from the existing $Q^{(\zeta)}_{D}(t)$. In case of a QD event, we need to sample the queue size and width from the unmodelled deeper part of the LOB. An important thing to note here is that sampling queue sizes depends on the queue width variable since a fatter queue will be expected to have a higher size than a thinner queue. Therefore we introduce this notation for the distribution of unseen queue sizes : $\Pi_{Q}(\hat{Q}_{D},m^{(\zeta)}_{D}(t^{+}))$. This has two parameters: first being $\hat{Q}_{D}$ which is calibrated according to the empirical data and the second being the width of the new fat-queue $m^{(\zeta)}_{D}(t^{+})$.
    If $m^{(\zeta)}_{D}(t)+m^{(\zeta)}_{T}(t)+\frac{1}{2}s(t)\leq M_{\frac{1}{2}}$, i.e. the width of the current deep wide queue is not violating the constraint in Defintion 2, we absorb the now empty levels in $m^{(\zeta)}_{T}(t)$ and create a new deep wide queue with the following width and volume:
@@
    $m^{(\zeta)}_{D}(t^{+})$ $=M_{\frac{1}{2}}-m^{(\zeta)}_{T}(t)-\frac{1}{2}s(t)$ (37)
    $Q^{(\zeta)}_{D}(t^{+})$ $\sim\Pi_{Q}(\hat{Q}_{D},m^{(\zeta)}_{D}(t^{+}))$ (38)
    If the constraint is violated, we set $m^{(\zeta)}_{D}(t^{+})$ to 1, sample $Q^{(\zeta)}_{D}(t^{+})$ from $\Pi_{Q}(\hat{Q}_{D},1)$ and let the deeper queue have just one non empty level since we cannot go deeper than $M_{\frac{1}{2}}+1$ in total depth of the LOB.
In summary, each of these 12 events and their effects on the coarse-grained LOB can be described, mathematically, by jump processes. Now we model the 12 events to form a mutually exciting 12 dimensional Hawkes Process with their intensities as follows. Let a $d$-dimensional Hawkes process have as the intensity of the process $\lambda^{(i)}(t)$ and the associated counting process $N^{(i)}(t)$ for $i=1,\ldots,d$. Here, $\mu^{(i)}$ is the exogenous intensity of the $i$-th dimension and $\phi^{(j\rightarrow i)}(t)$ is the excitation term from $j$-th dimension to $i$-th dimension. As described in [jain2023hawkes], we formulate the in-spread events’ intensity with a power-law function multiplier over the current spread in ticks with $\alpha$ (in \$) is defined as the scale parameter and $\beta$ is defined as the shape parameter. Since we do not constrain the excitations to be solely positive, in order for the intensities to make mathematical sense, we floor the intensities of each $i$ to zero. we For $d=12$, we have the following, with $s(t)$ being the spread-in-ticks at time $t$ and $\delta$ being the tick-size:
@@
3.3 Some Salient Features of the Model:
In this section we describe the dynamics of the some of the statistical features of a simulated LOB as defined above in Section 3 for empirical LOBs from the above model formulation.
3.3.1 Bid-Ask Spread and Mid-Price Changes:
It is easy to see that the spread process is decreased by $\eta_{IS}$ whenever an in-spread LO occurs and increased by $m^{(\zeta)}_{T}(t)$ whenever a queue depletion event (QD) occurs at the top queue on side $\zeta$. Therefore we can write the differential equation of spread as: